OpenStudy (anonymous):

Complete these ordered pairs for this equation. (0, ), (2, ), ( ,3) 3x - 2y = 6

OpenStudy (anonymous):

Substitute the corresponding values into the equation and solve for the other variable

OpenStudy (anonymous):

First example 3x - 2y = 6 x = 0 3(0) - 2y = 6 2y = -6 y = -3 (0 , -3)

OpenStudy (anonymous):

wich is the next one for 2, and what

OpenStudy (anonymous):

next is x = 2 Finally y = 3

OpenStudy (anonymous):

and then ,3 what is this one

OpenStudy (anonymous):

the ordered pairs are known as (x, y) and since you have the x in each ordered pair, just simply plug in each number for x and solve for the y

OpenStudy (anonymous):

i dont get it ßαbчღ

OpenStudy (anonymous):

The person above gaved you the answer

OpenStudy (anonymous):

but what is the other one

OpenStudy (anonymous):

What other one ?

OpenStudy (anonymous):

,3

OpenStudy (anonymous):

your answer is 0.3 ..

OpenStudy (anonymous):

3x - 2y = 6 The first ordered pair (0 , ). So x = 0. 3(0) - 2y = 6 2y = -6 y = -3 The first ordered pair is (0 , -3). Second ordered pair (2 , ). x = 2 3(2) - 2y = 6 6 - 2y = 6 2y = 0 y = 0 Second ordered pair is (2 , 0). Third ordered pair ( , 3). y = 3 3x - 2(3) = 6 3x - 6 = 6 3x = 12 x = 4 Third ordered pair is (4 , 3)

OpenStudy (anonymous):

did you just copy that @TURITW

OpenStudy (anonymous):

You mean I copied the work from somewhere or what? If that's the case, then no, i typed it out.

OpenStudy (anonymous):

Yeah that's what I ment and thats's nice Well love seems like Turitw here gaved you the answer and explained it step by step

OpenStudy (anonymous):

yea turitw

OpenStudy (anonymous):

i think those answers are wrong ßαbчღ or what

OpenStudy (anonymous):

They look correct to me sweety .

OpenStudy (anonymous):

but i put them and it didnt get it right sweety